# Exploiting Agent Symmetries for Performance Analysis of Distributed Optimization Methods

### v1

@@ 1 Introduction
[…]
@@
$x_{i}^{k+1}$ $=y_{i}^{k}-\alpha^{k}\nabla f_{i}(x_{i}^{k}),\hskip 56.9055pt\text{for $i=1,\dots,n$,}\vspace{-0.5mm}$ (3)
@@
This paper focuses on symmetric averaging matrices, but our results could be extended to non-symmetric ones. There are many other methods combining gradient sampling and consensus steps. We call $\mathcal{A}_{D}$ the set of all these decentralized optimization algorithms.
@@ 1.1 Contributions and paper organization
In [5], we have observed that for many performance settings, the worst-case guarantees obtained with PEP are independent of the number of agents, while the PEP problems are not. This motivates us to find compact ways of formulating the agent-dependent PEP. We have performed a first attempt in [4], where we built a relaxation of the problem whose size is independent of the number of agents $n$. In particular, the relaxation does not exploit separability of the objective function $f(x)$ in (1). While the relaxation in [4] gives good worst-case bounds, it remains an open question whether the equivalence with the agent-dependent formulation holds and how to interpret the resulting worst-case solution for the decentralized problem. In this paper, we provide an intuitive and systematic way of exploiting the agent symmetries in PEP to make the problem compact, with a size independent of the number of agents. In Section 3, we define equivalent agents in a PEP for decentralized optimization, and we leverage the convexity of the PEP to show that we can restrict it to solutions symmetrized over equivalent agents, without impacting its worst-case value. Section 4 focuses on the case where all the agents are equivalent, meaning none of them play a specific role in the algorithm or the performance evaluation. In that case, we show that the agent-dependent PEP can be written compactly, i.e. with an SDP whose size is independent of $n$. Moreover, we show that the worst-case value of this compact formulation is also independent of $n$ in many common performance settings of distributed optimization which are scale-invariant. This result is particularly powerful since it allows determining situations where the performance analysis of a distributed algorithm can be reduced to the fundamental case with only two agents. We further leverage agent equivalence to draw general conclusions about the symmetry of worst-case local functions and iterates in decentralized algorithms.
Then, Section 5 generalizes the results from Section 4 to situations where there are several equivalence classes of agents in the PEP, leading to a compact PEP formulation whose size only depends on the number of equivalence classes $m$, and not directly on the total number of agents $n$. Indeed, while many performance estimation problems for decentralized optimization have all agents equivalent, there are more complex ones that involve multiple equivalence classes of agents, for example, when different groups of agents use different (uncoordinated) step-sizes, function classes, initial conditions, or even different algorithms. It can also happen that the performance measure focuses on a specific group of agents, e.g. the performance of the worst agent. Efficiently and accurately assessing distributed optimization algorithm performance in such advanced settings would enhance comprehensive analysis and deepen our understanding of their behavior.
@@ 1.2 Related work
An alternative approach for automatic computation of performance guarantees of optimization method is proposed in [12] and relies on the formulation of optimization algorithms as dynamical systems. Integral quadratic constraints (IQC), generally used to obtain stability guarantees on complex dynamical systems, are adapted to provide sufficient conditions for the convergence of optimization methods and deduce numerical bounds on the convergence rates. This theory has been extended to decentralized optimization in [28]. The methodology analyzes the convergence rate of a single iterate, which is beneficial for the problem size that remains small, and that is also independent of the number of agents in the decentralized case. However, this does not allow dealing with non-geometric convergences, e.g. on smooth convex functions, nor to analyze cases where a property applies over several iterations, e.g. when averaging matrices are constant. These are possible with the PEP approach which computes the worst-case performance on a given number $t$ of iterations, but solving problems whose size grows with $t$. Moreover, the current IQC approach only applies to settings where all the agents are equivalent, which is not the case, for example, if we want to analyze the performance of the worst agent. In this work, we propose a practical way, via the PEP approach, to compute performance where there are several equivalence classes of agents. This expands the range of situations we can automatically analyze efficiently in decentralized optimization, including more complex or realistic settings.
@@ 2.1 Performance Estimation Problem (PEP) framework
$w(\mathcal{S}_{n})\leavevmode\nobreak\ =\leavevmode\nobreak$ $\underset{x^{*},\{x_{i}^{k},y_{i}^{k},f_{i}\}_{i\in\mathcal{V}}}{\max}\leavevmode\nobreak\ \mathcal{P}(f_{i},x_{i}^{0},\dots,x_{i}^{t},x^{*})$ =$\stackrel{{\scriptstyle\text{e.g.}}}{{=}}\leavevmode\nobreak\ \text{\small{$\frac{1}{n}\sum_{i=1}^{n}\quantity(f_{i}(\overline{x}^{t})-f_{i}(x^{*}))$}}$ (7)
@@
$f_{i}\in\mathcal{F},$ for $\text{for }\leavevmode\nobreak\ i\in\mathcal{V}$ (class of functions) (9)
[…]
@@
We assume here that the local functions all belong to the same function class. This assumption is usually done in the literature but is not necessary in this PEP formulation. The problem can also have other auxiliary iterates as variables if needed for the analyzed algorithm. Here we have chosen to show auxiliary iterates $y_{i}^{k}$ which are the results of the consensus steps (4) that may be used at each iteration of the algorithm.
The optimal value of problem (7), denoted $w(\mathcal{S}_{n})$, gives by definition a tight worst-case performance bound for the given setting $\mathcal{S}_{n}$. Moreover, the optimal solution corresponds to an instance of local functions and initial points actually reaching this upper bound, which can provide very relevant information on the bottlenecks faced by the algorithm. Solving a PEP such as (7) is in general not easy because the problem is inherently infinite-dimensional, as it contains continuous functions $f_{i}$ among its variables. Nevertheless, Taylor et al. have shown [30, 29] that PEPs can be formulated as a finite semidefinite program (SDP) and can thus be solved exactly, for a wide class of centralized first-order algorithms and different classes of functions. The reformulation techniques developed for classical optimization algorithms can also be applied for distributed optimization, as detailed in [5]. In what follows, we briefly explain how to reformulate the problem (7) into an SDP. One of the differences in PEP for distributed optimization is that the problem must find the worst-case for each of the $n$ local functions $f_{i}$ and the $n$ sequences of local iterates $x_{i}^{0}\dots x_{i}^{t}$ ($i=1,\dots,n$). To render problem (7) finite, for each agent $i$, rather than considering its local function $f_{i}$ as a whole, we only consider the discrete set $\{\quantity(x_{i}^{k},g_{i}^{k},f_{i}^{k})\}_{k\in I=\{0,\dots,t,*\}}$ of local iterates $x_{i}^{k}$ together with their local gradient-vectors $g_{i}^{k}$ and local function values $f_{i}^{k}$. We then impose interpolation constraints on this set to ensure its consistency with an actual function $f_{i}\in\mathcal{F}$, in the sense that the set is $\mathcal{F}$-interpolable: there exists a function $f_{i}\in\mathcal{F}$ such that $f_{i}(x_{i}^{k})=f_{i}^{k}$ and $\nabla f_{i}(x_{i}^{k})=g_{i}^{k}$. Such interpolation constraints are provided for many classes of functions in [29, Section 3], such as the class $L$-smooth and $\mu$-strongly convex functions.
@@
$f_{k}-f_{l}-g_{l}^{T}(x_{k}-x_{l})\geq\frac{1}{2(1-\mu/L)}\quantity(\frac{1}{L}\|g_{k}-g_{l}\|^{2}+\mu\|x_{k}-x_{l}\|^{2}-2\frac{\mu}{L}(g_{k}-g_{l})^{T}(x_{k}-x_{l})).$
Then, all the constraints from (7) can be expressed in terms of $\{y_{i}^{k},x_{i}^{k},g_{i}^{k},f_{i}^{k}\}_{k\in I=\{0,\dots,t,*\}}$. When the averaging matrix $W$ is given, i.e. $M=\{W\}$, the algorithm constraints simply consist of linear constraints between these variables. The optimality condition for (1) can be expressed as a linear constraint on the local gradients at $x^{*}$. The performance criterion, the initial conditions, and the interpolation constraints are usually quadratic and potentially non-convex in the local iterates and local gradients but they are linear in the scalar products of these and in the function values. Hence, letting the decision variables of the PEP be these scalar products and the function values, one can reformulate it as a semi-definite program (SDP) which can be solved efficiently. For this purpose, we define a vector of functions values $F$ and a Gram matrix $G$ that contains scalar products between all vectors, e.g. the local iterates $y_{i}^{k},x_{i}^{k}\in\mathbb{R}^{d}$ and the local gradients vectors $g_{i}^{k}\in\mathbb{R}^{d}$.
@@
By definition, $G$ is symmetric and positive semidefinite. Moreover, every matrix $G\succeq 0$ corresponds to a matrix $P$ whose number of rows is equal to $\text{rank}\leavevmode\nobreak\ G$. It has been shown in [29] that the reformulation of a PEP, such as (7), with $G$ as variable is lossless provided that we use necessary and sufficient interpolation constraints for $\mathcal{F}$ and that we do not impose the dimension $d$ of the worst-case, and indeed look for the worst-case over all possible dimensions (imposing the dimension would correspond to adding a typically less tractable rank constraint on $G$). The resulting SDP-PEP takes thus the form of
$w(\mathcal{S}_{n})\quad=\hskip 28.45274pt$ $\underset{F,G}{\mathrm{max}}\quad\mathcal{P}(F,G)$ (14)
[…]
$\quad F,G\quad\text{satisfy}$ algorithm constraints for $\mathcal{A}$ (16)
interpolation constraints for ,$\text{interpolation constraints for $\mathcal{F}$},$ (17)
initial conditions ,$\text{initial conditions $\mathcal{I}$},$ (18)
[…]
Such an SDP formulation is convenient because it can be solved numerically to global optimality, leading to a tight worst-case bound, and it also provides the worst-case solution over all possible problem dimensions. We refer the reader to [29] for more details about the SDP formulation of a PEP, including ways of reducing the size of matrix $G$. However, the dimension of $G$ always depends on the number of iterations $t$, and on the number of agents $n$.
From a solution $G$, $F$ of the SDP formulation, we can construct a solution for the variables $\{y_{i}^{k},x_{i}^{k},g_{i}^{k},f_{i}^{k}\}_{i\in\mathcal{V},k\in I}$, e.g. using the Cholesky decomposition of $G$. For each agent $i\in\mathcal{V}$, its resulting set of points satisfies the interpolation constraints, so we can construct its corresponding worst-case local function from $\mathcal{F}$ interpolating these points. Proposition 2 below states sufficient conditions under which a PEP on a distributed optimization method can be formulated as an SDP. These conditions are satisfied in many PEP settings, allowing to formulate and solve the SDP PEP formulation for any distributed first-order methods from $\mathcal{A}_{D}$, with many classes of local functions, initial conditions, and performance criteria, see [29], [5]. The proposition uses the following definition.
@@
Consider a Gram matrix $G$ and a vector $F$, as defined in (12). We say that an expression, such as a constraint or an objective, is linearly (resp. LMI) Gram-representable if it can be expressed using a finite set of linear (resp. LMI) constraints or expressions involving (part of) $G$ and $F$.
Proposition 2 ([29, Proposition 2.6]).
Let $w(\mathcal{S}_{n}=\{n,\mathcal{A},t,\mathcal{P},\mathcal{F},\mathcal{I},\mathcal{W}\})$ be the worst-case performance of the execution of $t$ iterations of distributed algorithm $\mathcal{A}$ with $n\geq 2$ agents, with respect to the performance criterion $\mathcal{P}$, valid for any starting point satisfying the set of initial conditions $\mathcal{I}$, any local functions in a given set of functions $\mathcal{F}$ and any averaging matrix in the set of matrices $\mathcal{W}$. If $\mathcal{A},\mathcal{P},\mathcal{I}$ and interpolation constraints for $\mathcal{F}$ and for $\mathcal{W}$ are linearly (or LMI) Gram representable, then the computation of $w(\mathcal{S}_{n})$ can be formulated as an SDP, with $G\succeq 0$ and $F$ as variables, see (14).
@@
In [29], Definition 3 and Proposition 2 were only formulated for linearly Gram-representable constraints, but their extension to LMI Gram-representable constraints are direct and were already introduced in [5] in which LMI constraints are used to represent consensus steps in PEP.
As briefly mentioned, when the set of possible averaging matrices $\mathcal{W}$ contains only one given matrix $W$, i.e. we know the averaging matrix used, the PEP framework presented here is complete and provides worst-case bounds specific to this value of $W$. However, the literature on distributed optimization generally provides bounds on the performance of an algorithm that are valid over a larger set of averaging matrices $\mathcal{W}$, typically based on its spectral properties, see [20] for a survey. Such bounds allow better characterization of the general behavior of the algorithm and can be applied in a wider range of settings. Obtaining such bound with PEP requires to have a tractable representation of the set of all pairs of variables $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}$ that can be involved in consensus steps with the same matrix from the given matrix class. Our previous work [5] proposes and uses necessary constraints for describing the set of symmetric and doubly-stochastic matrices with a given range on the non-principal eigenvalues. The next section describes these constraints and analyzes their sufficiency.

### v2

@@ 1 Introduction
$\underset{\text{\normalsize$x\in\mathbb{R}^{d}$}}{\mathrm{min}}\quad f(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x),\vskip-2.84526pt$ (1)
@@
$x_{i}^{k+1}$ $=y_{i}^{k}-\alpha^{k}\nabla f_{i}(x_{i}^{k}),\hskip 56.9055pt\text{for $i=1,\dots,n$,}\vskip-1.42262pt$ (3)
@@
There are many other methods combining gradient sampling and consensus steps. We call $\mathcal{A}_{D}$ the set of all these decentralized optimization algorithms.
@@ 1.1 Contributions and paper organization
In [5], we have also observed that for many performance settings, the worst-case guarantees obtained with PEP are independent of the number of agents, while the PEP problems are not. This motivates us to find compact ways of formulating the agent-dependent PEP. We have performed a first attempt in [4], where we built a relaxation of the problem whose size is independent of the number of agents $n$. In particular, the relaxation does not exploit separability of the objective function $f(x)$ in (1). While the relaxation in [4] gives good worst-case bounds, it remains an open question whether the equivalence with the agent-dependent formulation holds and how to interpret the resulting worst-case solution for the decentralized problem. In this paper, we provide an intuitive and systematic way of exploiting the agent symmetries in PEP to make the problem compact, with a size independent of the number of agents. In Section 3, we define equivalent agents in a PEP for decentralized optimization, and we leverage the convexity of the PEP to show that we can restrict it to solutions symmetrized over equivalent agents, without impacting its worst-case value. Section 4 focuses on the case where all the agents are equivalent, meaning none of them play a specific role in the algorithm or the performance evaluation. In that case, we show that the agent-dependent PEP can be written compactly, i.e. with an SDP whose size is independent of $n$. Moreover, we show that the worst-case value of this compact formulation is also independent of $n$ in many common performance settings of distributed optimization which are scale-invariant. This result is particularly powerful since it allows determining situations where the performance analysis of a distributed algorithm can be reduced to the fundamental case with only two agents. We further leverage agent equivalence to draw general conclusions about the symmetry of worst-case local functions and iterates in decentralized algorithms.
Then, Section 5 generalizes the results from Section 4 to situations where there are several equivalence classes of agents in the PEP, leading to a compact PEP formulation whose size only depends on the number of equivalence classes $r$, and not directly on the total number of agents $n$. Indeed, while many performance estimation problems for decentralized optimization have all agents equivalent, there are more complex ones that involve multiple equivalence classes of agents, for example, when different groups of agents use different (uncoordinated) step-sizes, function classes, initial conditions, or even different algorithms. It can also happen that the performance measure focuses on a specific group of agents, e.g. the performance of the worst agent. Efficiently and accurately assessing distributed optimization algorithm performance in such advanced settings would enhance comprehensive analysis and deepen our understanding of their behavior.
@@ 1.2 Related work
An alternative approach for automatic computation of performance guarantees of optimization methods is proposed in [13] and relies on the formulation of optimization algorithms as dynamical systems. Integral quadratic constraints (IQC), generally used to obtain stability guarantees on complex dynamical systems, are adapted to provide sufficient conditions for the convergence of optimization methods and deduce numerical bounds on the convergence rates. This theory has been extended to decentralized optimization in [29]. The methodology analyzes the convergence rate of a single iterate, which is beneficial for the problem size that remains small, and that is also independent of the number of agents in the decentralized case. However, this does not allow dealing with non-geometric convergences, e.g. on smooth convex functions, nor to analyze cases where a property applies over several iterations, e.g. when averaging matrices are constant. These are possible with the PEP approach which computes the worst-case performance on a given number $t$ of iterations, but solving problems whose size grows with $t$. Moreover, the current IQC approach only applies to settings where all the agents are equivalent, which is not the case, for example, if we want to analyze the performance of the worst agent. In this work, we propose a practical way, via the PEP approach, to compute performance where there are several equivalence classes of agents. This expands the range of situations we can automatically analyze efficiently in decentralized optimization, including more complex or realistic settings.
@@ 2.1 Performance Estimation Problem (PEP) framework
$w(\mathcal{S}_{n})\penalty 10000\ =\penalty 10000\$$\underset{x^{*},\{x_{i}^{k},y_{i}^{k},f_{i}\}_{i\in\mathcal{V}}}{\max}\penalty 10000\ \mathcal{P}(f_{i},x_{i}^{0},\dots,x_{i}^{t},x^{*})$ $\stackrel{{\scriptstyle\text{e.g.}}}{{=}}\penalty 10000\ \text{\small{$\frac{1}{n}\sum_{i=1}^{n}\quantity(f_{i}(\overline{x}^{t})-f_{i}(x^{*}))$}}$ (7)
@@
$f_{i}\in\mathcal{F},$ $\text{for }\penalty 10000\ i\in\mathcal{V}$ (class of functions) (9)
$x_{i}^{0}\quad\underset{\text{e.g. $\|x_{i}^{0}-x^{*}\|^{2}\leq 1$,\penalty 10000\ $i\in\mathcal{V}$}}{\text{satisfies $\mathcal{I}$, \hskip 56.9055pt}}$ (initial conditions) (10)
@@
We assume here that the local functions all belong to the same function class. This assumption is usually made in the literature but is not necessary in this PEP formulation. The problem can also have other auxiliary iterates as variables if needed for the analyzed algorithm. Here we have chosen to show auxiliary iterates $y_{i}^{k}$ which are the results of the consensus steps (4) that may be used at each iteration of the algorithm.
The optimal value of problem (7), denoted $w(\mathcal{S}_{n})$, gives by definition a tight worst-case performance bound for the given setting $\mathcal{S}_{n}$. Moreover, the optimal solution corresponds to an instance of local functions and initial points actually reaching this upper bound, which can provide very relevant information on the bottlenecks faced by the algorithm. Solving a PEP such as (7) is in general not easy because the problem is inherently infinite-dimensional, as it contains continuous functions $f_{i}$ among its variables. Nevertheless, Taylor et al. have shown [32, 30] that PEPs can be formulated as a finite semidefinite program (SDP) and can thus be solved exactly, for a wide class of centralized first-order algorithms and different classes of functions. The reformulation techniques developed for classical optimization algorithms can also be applied for distributed optimization, as detailed in [5]. In what follows, we briefly explain how to reformulate the problem (7) into an SDP. One of the differences in PEP for distributed optimization is that the problem must find the worst-case for each of the $n$ local functions $f_{i}$ and the $n$ sequences of local iterates $x_{i}^{0}\dots x_{i}^{t}$ ($i=1,\dots,n$). To render problem (7) finite, for each agent $i$, rather than considering its local function $f_{i}$ as a whole, we only consider the discrete set $\{\quantity(x_{i}^{k},g_{i}^{k},f_{i}^{k})\}_{k\in I=\{0,\dots,t,*\}}$ of local iterates $x_{i}^{k}$ together with their local gradient-vectors $g_{i}^{k}$ and local function values $f_{i}^{k}$. We then impose interpolation constraints on this set to ensure its consistency with an actual function $f_{i}\in\mathcal{F}$, in the sense that the set is $\mathcal{F}$-interpolable : there exists a function $f_{i}\in\mathcal{F}$ such that $f_{i}(x_{i}^{k})=f_{i}^{k}$ and $\nabla f_{i}(x_{i}^{k})=g_{i}^{k}$. Such interpolation constraints are provided for many classes of functions in [30, Section 3], such as the class $L$-smooth and $\mu$-strongly convex functions.
@@
$f_{k}-f_{l}-g_{l}^{T}(x_{k}-x_{l})\geq\frac{1}{2(1-\mu/L)}\quantity(\frac{1}{L}\|g_{k}-g_{l}\|^{2}+\mu\|x_{k}-x_{l}\|^{2}-2\frac{\mu}{L}(g_{k}-g_{l})^{T}(x_{k}-x_{l})).$ (12)
Then, all the constraints from (7) can be expressed in terms of $\{y_{i}^{k},x_{i}^{k},g_{i}^{k},f_{i}^{k}\}_{k\in I=\{0,\dots,t,*\}}$. When the averaging matrix $W$ is given, i.e. $\mathcal{W}=\{W\}$, the algorithm constraints simply consist of linear constraints between these variables. The optimality condition for (1) can be expressed as a linear constraint on the local gradients at $x^{*}$. The performance criterion, the initial conditions, and the interpolation constraints are usually quadratic and potentially non-convex in the local iterates and local gradients but they are often linear in the scalar products of these and in the function values, see (12) for example. In such cases, letting the decision variables of the PEP be these scalar products and the function values, one can reformulate it as a semi-definite program (SDP) which can be solved efficiently. For this purpose, we define a vector of function values $\mathbf{f}$ and a Gram matrix $G$ that contains scalar products between all vectors, e.g. the local iterates $y_{i}^{k},x_{i}^{k}\in\mathbb{R}^{d}$ and the local gradient vectors $g_{i}^{k}\in\mathbb{R}^{d}$.
$\mathbf{f}=[f_{i}^{k}]_{i\in\mathcal{V},\penalty 10000\ k\in I=\{0,\dots,t,*\}}\penalty 10000\ \in\mathbb{R}^{n(t+2)},\hskip 34.14322ptG=P^{T}P\penalty 10000\ \in\mathbb{R}^{3n(t+2)\times 3n(t+2)},\text{ with }P=\quantity[y_{i}^{k}\penalty 10000\ x_{i}^{k}\penalty 10000\ g_{i}^{k}]_{i\in\mathcal{V},k\in I}.$ (13)
By definition, $G$ is symmetric and positive semidefinite. Moreover, every matrix $G\succeq 0$ corresponds to a matrix $P$ whose number of rows is equal to $\text{rank}\penalty 10000\ G$. It has been shown in [30] that the reformulation of a PEP, such as (7), with $G$ as variable is lossless provided that we use necessary and sufficient interpolation constraints for $\mathcal{F}$ and that we do not impose the dimension $d$ of the worst-case, and indeed look for the worst-case over all possible dimensions (imposing the dimension would correspond to adding a typically less tractable rank constraint on $G$). The resulting SDP-PEP thus takes the form of
$w(\mathcal{S}_{n})\quad=\hskip 28.45274pt$ $\underset{\mathbf{f},G}{\mathrm{max}}\quad\mathcal{P}(\mathbf{f},G)$ (15)
$\text{ s.t.}\qquad G\succeq 0,$ (16)
$\quad\mathbf{f},G\quad\text{satisfy}$ algorithm constraints for $\mathcal{A}$ (17)
interpolation constraints for $\hskip-142.26378pt\text${interpolation constraints for $\mathcal{F}$$}, (18)
initial conditions $\hskip-142.26378pt\text${initial conditions $\mathcal{I}$$}, (19)
@@
where the objective function $\mathcal{P}$ and the constraints are linear in $\mathbf{f}$ and $G$. The performance criterion function $\mathcal{P}$ in (15) corresponds to the same function as the function $\mathcal{P}$ in (7), up to a change of variables. We slightly abuse notations to denote these two functions with the same symbol $\mathcal{P}$, for readibility.
Such an SDP formulation (15) is convenient because it can be solved numerically to global optimality, leading to a tight worst-case bound, and it also provides the worst-case solution over all possible problem dimensions. We refer the reader to [30] for more details about the SDP formulation of a PEP, including ways of reducing the size of matrix $G$. However, the dimension of $G$ always depends on the number of iterations $t$, and on the number of agents $n$.
From a solution $G$, $\mathbf{f}$ of the SDP formulation, we can construct a solution for the variables $\{y_{i}^{k},x_{i}^{k},g_{i}^{k},f_{i}^{k}\}_{i\in\mathcal{V},k\in I}$, e.g. using the Cholesky decomposition of $G$. For each agent $i\in\mathcal{V}$, its resulting set of points satisfies the interpolation constraints, so we can construct its corresponding worst-case local function from $\mathcal{F}$ interpolating these points. Proposition 2 below states sufficient conditions under which a PEP on a distributed optimization method can be formulated as an SDP. These conditions are satisfied in many PEP settings, allowing to formulate and solve the SDP PEP formulation for any distributed first-order methods from $\mathcal{A}_{D}$, with many classes of local functions, initial conditions, and performance criteria, see [30, 5, 6]. The proposition uses the following definition.
@@
Consider a Gram matrix $G$ and a vector $\mathbf{f}$, as defined in (13). We say that an expression, such as a constraint or an objective, is linearly (resp. LMI) Gram-representable if it can be expressed using a finite set of linear (resp. LMI) constraints or expressions involving (part of) $G$ and $\mathbf{f}$.
Proposition 2 ([30, Proposition 2.6]).
Let $w(\mathcal{S}_{n}=\{n,\mathcal{A},t,\mathcal{P},\mathcal{F},\mathcal{I},\mathcal{W}\})$ be the worst-case performance of the execution of $t$ iterations of distributed algorithm $\mathcal{A}$ with $n\geq 2$ agents, with respect to the performance criterion $\mathcal{P}$, valid for any starting point satisfying the set of initial conditions $\mathcal{I}$, any local functions in a given set of functions $\mathcal{F}$ and any averaging matrix in the set of matrices $\mathcal{W}$. If $\mathcal{A},\mathcal{P},\mathcal{I}$ and interpolation constraints for $\mathcal{F}$ and for $\mathcal{W}$ are linearly (or LMI) Gram representable, then the computation of $w(\mathcal{S}_{n})$ can be formulated as an SDP, with $G\succeq 0$ and $\mathbf{f}$ as variables, see (15).
@@
In [30], Definition 3 and Proposition 2 were only formulated for linearly Gram-representable constraints, but their extensions to LMI Gram-representable constraints are direct and were already introduced in [5] in which LMI constraints are used to represent consensus steps in PEP.
As briefly mentioned, when the set of possible averaging matrices $\mathcal{W}$ contains only one given matrix $W$, i.e. we know the averaging matrix used, the PEP framework presented here is complete and provides worst-case bounds specific to this value of $W$. However, the literature on distributed optimization generally provides bounds on the performance of an algorithm that are valid over a larger set of averaging matrices $\mathcal{W}$, typically based on its spectral properties, see [21] for a survey. Such bounds allow better characterization of the general behavior of the algorithm and can be applied in a wider range of settings. Obtaining such bounds with PEP requires having a tractable representation of the set of all pairs of variables $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}$ that can be involved in consensus steps with the same matrix from the given matrix class. Our previous work [5] proposes and uses necessary constraints for describing the set of symmetric and doubly-stochastic matrices with a given range on the non-principal eigenvalues. The next section describes these constraints and analyzes their sufficiency.